OpenStudy (anonymous):

tan theta + sec theta = 1 between 0 and 2 pi

OpenStudy (swissgirl):

can theta=0 ? Like is is\[ 0< \theta <\Pi\] Or \[ 0 \leq \theta \leq \Pi\] ?

OpenStudy (anonymous):

0<θ<2Π has to be between this interval

OpenStudy (anonymous):

use substitution y=tan(theta/2)